OpenStudy (anonymous):

PLEASE HELP im so confused write an equation in point-slope form of the line through point J(4, 1) with slope -4

OpenStudy (anonymous):

I don't know how to do this

OpenStudy (anonymous):

Point slope equation: y - y1 = m(x - x1) (x1, y1) is any point on the line. You know a point on your line, so plug that in! m is the slope. You know that too, so plug it in.

OpenStudy (jdoe0001):

\(\large {\textit{point-slope form}\implies y-y_1=m(x-x_1)\\ \quad \\ slope=m=-4\qquad \begin{array}{lllll} &x_1&y_1\\ &(4\quad ,&1) \end{array}}\)

OpenStudy (anonymous):

y-y1=m(x-x1) (x1,y1)=(4,1) m=-4

OpenStudy (anonymous):

woah I get it now thank you id give all 3 of u guys medals but it prob wont let me